OpenStudy (anonymous):

express the radical in simplest form. square root of 4/6

OpenStudy (amoodarya):

\[\sqrt{\frac{ 4 }{ 6 }}\]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (amoodarya):

\[\sqrt{\frac{ 4 }{ 6 }}=\\ \sqrt{\frac{ 2 }{ 3 }}=\\\ \sqrt{\frac{ 2 }{ 3 }\frac{ 3 }{ 3 }}=\\\sqrt{\frac{ 6 }{ 9 }}=\\ \frac{\sqrt{6}}{\sqrt{9}}=\\ \frac{ \sqrt{6}}{ 3}\]

OpenStudy (anonymous):

thanks can you help me with a few more?